A 4.80 g lead bullet traveling at 500 m/s is stopped by a large tree. If half the kinetic energy of the bullet is transformed into internal energy and remains with the bullet while the other half is transmitted to the tree, what is the increase in temperature of the bullet?
Answer #1

As we know that;

The Kinetic Energy is,

= K.E= 0.5*m*v^2

So,

= K.E= 0.5*4.8*10^-3*500^2

= K.E = 600 J

Then,

=Q = msdT

= 0.5*600 = 5*0.128*dT

= 300=0.64*dT

=dT=300/0.64

= dT = 468.7°C